# A Practical Adaptive Subgame Perfect Gradient Method

### v1

@@ A.1 Preconditioning Algorithms
We briefly present standard algorithms for efficient computation of matrix-vector products with the inverse Hessian approximation $B$ and the Hessian approximation $B^{-1}$. Algorithm˜6 follows the standard two-loop recursion method [1]. Algorithm˜7 simply uses the compact representation [50] of the Hessian approximation $B^{-1}$,
@@
$S=\begin{bmatrix}s_{1}&\dots&s_{t}\end{bmatrix}\quad Y=\begin{bmatrix}y_{1}&\dots&y_{t}\end{bmatrix}$
@@
Thus, to demonstrate the minimax optimality we have to provide a matching lower bound. To do so, we will construct a particular oracle that is hard for all $A\in\mathcal{A}^{\mathcal{H}}$.
@@
Next, we introduce some notation. Let $\phi_{n}$ be the solution to (3.12) at iteration $n$, (i.e., $\phi_{n}=\sum_{i={n-k}}^{n-1}{\rho}_{i}{\tau}_{i}+\sum_{i={n-k}}^{n-1}{\gamma}_{i}$) and define $\tau_{n,i}$ as in Theorem˜3.3. Further define
@@
We now consider the dual program to (A.2) and obtain the following lemma. Recall from Lemma˜3.1, that a bounded solution to (3.12) implies that strong duality holds. The derivation of the dual problem below is effectively identical to that in [5, Lemma 3].
@@
$\inf_{\xi,y}\left\{\begin{array}[]{ll}\frac{2\|y\|^{2}}{\xi L}+\frac{\xi}{2}\|x_{0}\|^{2}+2\langle x_{0},y\rangle:&\begin{array}[]{l}\xi(-h+v_{m}\tau)-2Z^{T}y\geq\tau\\ \xi(-w+v_{m}\mathbf{1})+2G^{T}y\geq 1\\ \xi>0\end{array}\end{array}\right\}$ (A.3)
@@
We continue following the approach of [5] and define $f_{\star}=v_{m}-\frac{1}{\xi}$. Then our feasibility constraints for $\xi$ can be rewritten as
$\tau_{i}\left(f_{\star}-f_{i}+\frac{1}{2L}\|g_{i}\|^{2}\right)$ $+\frac{L}{2}\|x_{0}\|^{2}-\frac{L}{2}\|z_{i+1}\|^{2}+L\langle z_{i+1}-x_{0},z^{\prime}\rangle\geq 0\qquad\forall i\in[n-k,n-1]$ (A.4)
$f_{\star}-f_{i}-\langle g_{i},z^{\prime}-x_{i}\rangle\geq 0\qquad\forall i\in[n-k,n-1].$ (A.5)
@@
where $\beta_{i}$ is defined inductively as $\beta_{i}=\frac{1}{\phi_{i}(2\psi_{i}+1)}\left(1+\sum_{j=n}^{i-1}\psi_{j}^{2}\beta_{j}\right)$ with $\beta_{n}=\frac{1}{\phi_{n}(2\psi_{n}+1)}$. Note that this oracle selection is independent of the value of $x_{i}$; this is reasonable since our oracle is not restricted to return values corresponding to an actual function [^6] Indeed, this construction even allows the oracle to be “inconsistent” over time: If an algorithm selects $x_{1}=x_{0}$, the oracle can return $(f_{1},g_{1})$ with $f_{1}\neq f_{0}$ and $g_{1}\neq g_{0}$, as long as $Q_{0,1}(L)\geq 0$ still holds.. It is only required to satisfy $\{Q_{i,j}\}_{i<j<n}\cup\{W_{\star,i}\}_{i=0}^{n-1}$. Lastly, we set $x_{\star}=z_{N+1}$.
@@
and the result follows from induction.
@@ A.3 Additional Numerical Results
For reference, we include performance plots for each individual problem type considered in Section˜4.2.

### v2

@@ A.1 Preconditioning Algorithms
We briefly present standard algorithms for efficient computation of matrix-vector products with the inverse Hessian approximation $B$ and the Hessian approximation $B^{-1}$. Algorithm˜6 follows the standard two-loop recursion method [43]. Algorithm˜7 uses the compact representation [7] of the Hessian approximation $B^{-1}$,
@@
$S=\begin{bmatrix}s_{1}&\cdots&s_{t}\end{bmatrix}\quad Y=\begin{bmatrix}y_{1}&\cdots&y_{t}\end{bmatrix}$
@@
Thus, to demonstrate the minimax optimality, it suffices to provide a matching lower bound. To do so, we will construct a particular oracle that is hard for all $A\in\mathcal{A}^{\mathcal{H}}$.
@@
Next, we introduce some notation. Let $\phi_{n}$ be the optimal value of (3.12) at iteration $n$, (i.e., $\phi_{n}=\sum_{i={n-k}}^{n-1}{\rho}_{i}{\tau}_{i}+\sum_{i={n-k}}^{n-1}{\gamma}_{i}$) and define $\tau_{n,i}$ as in Theorem˜3.3. Further define
@@
We now consider the dual program to (A.2) and obtain the following lemma. Recall from Lemma˜3.1 that a bounded solution to (3.12) implies that strong duality holds. The derivation of the dual problem below is effectively identical to that in [19, Lemma 3].
@@
$\left\{\begin{array}[]{ll}\inf_{\xi,y}&\frac{2\|y\|^{2}}{\xi L}+\frac{\xi}{2}\|x_{0}\|^{2}+2\langle x_{0},y\rangle\\ \text{\emph{s.t.}}&\xi(-h+v_{m}\tau)-2Z^{T}y\geq\tau\\ &\xi(-w+v_{m}\mathbf{1})+2G^{T}y\geq 1\\ &\xi>0\end{array}\right$. (A.3)
@@
We continue following the approach of [19] and define $f_{\star}=v_{m}-\frac{1}{\xi}$. Then our feasibility constraints for $\xi$ can be rewritten for all $i\in[n-k,n-1]$ as
$\tau_{i}\left(f_{\star}-f_{i}+\frac{1}{2L}\|g_{i}\|^{2}\right)$ $+\frac{L}{2}\|x_{0}\|^{2}-\frac{L}{2}\|z_{i+1}\|^{2}+L\langle z_{i+1}-x_{0},z^{\prime}\rangle\geq 0$ (A.4)
$f_{\star}-f_{i}-\langle g_{i},z^{\prime}-x_{i}\rangle\geq 0.$ (A.5)
@@
where $\beta_{i}$ is defined inductively as $\beta_{i}=\frac{1}{\phi_{i}(2\psi_{i}+1)}\left(1+\sum_{j=n}^{i-1}\psi_{j}^{2}\beta_{j}\right)$ with $\beta_{n}=\frac{1}{\phi_{n}(2\psi_{n}+1)}$. Note that this oracle selection is independent of the value of $x_{i}$; this is permissible since our oracle is not restricted to return values corresponding to an actual function [^7] Indeed, this construction even allows the oracle to be “inconsistent” over time: If an algorithm selects $x_{1}=x_{0}$, the oracle can return $(f_{1},g_{1})$ with $f_{1}\neq f_{0}$ and $g_{1}\neq g_{0}$, as long as $Q_{0,1}(L)\geq 0$ still holds.. It is only required to satisfy $\{Q_{i,j}\}_{i<j<n}\cup\{W_{\star,i}\}_{i=0}^{n-1}$. Lastly, we set $x_{\star}=z_{N+1}$.
@@
and the result follows from induction. ∎
@@ A.3 Additional Numerical Results
For completeness, we include performance plots for each individual problem type considered in Section˜4.2.